OpenStudy (anonymous):

Choose the correct simplification of the expression −5x2(4x − 6x2 − 3). 30x4 − 20x3 + 15x2 −11x4 − x3 − 8x2 −30x4 + 20x3 − 15x2 30x4 + 20x3 + 15x2

OpenStudy (notamathgenius):

@Kitt020912 @amistre64

OpenStudy (espex):

Distribute your \(-5x^2\) into your parenthesis to solve this.

OpenStudy (anonymous):

i got -2x2 + 30x2+15x2

OpenStudy (espex):

How did you get that answer?

OpenStudy (anonymous):

by distribute -5x2

OpenStudy (espex):

\[-5x^2*4x=?\]\[-5x^2*-6x^2=?\]\[-5x^2*-3=?\]

OpenStudy (anonymous):

15x2

OpenStudy (espex):

For the last one, yes, \(15x^2\) is the answer.

OpenStudy (anonymous):

thanx

OpenStudy (espex):

You still have two more terms to complete the simplification.

OpenStudy (anonymous):

So it's all solved now? Do you need anymore help, @dreap3 c:

OpenStudy (anonymous):

ummm i think so

OpenStudy (anonymous):

Just close this question and open another c:

OpenStudy (espex):

For the record, @dreap3 did not present a correct answer and implied that \(15x^2\) might have been it when it was only the final term in the simplified expression.

OpenStudy (anonymous):

dont worry i solved it i got the answer